# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v1

@@ IV-A Analysis of the sufficient condition for $b$
IV-B Theoretical vs. simulated MSE of the sliding DFT unfolding
@@
where $p_{\mathrm{rc}}(t)$ is a raised-cosine pulse with roll-off parameter $\beta=1$. The peak amplitude of the raised-cosine pulse $p_{\mathrm{rc}}(t)$ is set to unity. The pulse amplitude $A_{m}$ is drawn uniformly from the discrete set $\{-2,-1,0,+1,+2\}$, and $T$ is the time difference between two consecutive pulses. Intuitively, the signal $f(t)$ is constructed by generating 100,000 raised-cosine pulses with random amplitudes at every symbol period $T$. This signal is fed to a low-resolution 4-bit modulo ADC with sampling rate $\frac{1}{T_{\mathrm{s}}}=\frac{\mathrm{OF}\omega_{m}}{2\pi}$ to produce the quantized modulo samples $f_{\lambda^{\prime},\mathrm{q}}[n]$. The dither sequence and modulo threshold are configured according to the modulo ADC parameters.
@@ V Summary and Future Directions
It is worth mentioning that while the proposed unfolding scheme relies on the availability of 1-bit folding information signal $c[n]$, the proposed recovery can be modified to work without $c[n]$. This can be done by performing support recovery on $\underline{\tilde{z}^{(i)}_{w}}[n]$. This idea, together with its performance guarantees, is considered a potential subject of future work. We also note that the sufficient conditions derived for the oversampling factor and ADC resolution are not tight. Hence, another future direction is to tighten these sufficient conditions. Lastly, it is also of interest to apply this proposed recovery algorithm to specific low-resolution signal processing systems (e.g., communication receivers, radars, etc.).
@@ Appendix A Proof of Proposition 1
$K=$ $N-2\left\lfloor\frac{\rho K}{2}\right\rfloor-K_{\mathrm{SL}}$
$\geq$ $N(1-\rho)-K_{\mathrm{SL}}.$ (23)
@@
Here, we used $\mathrm{OF}=\frac{1}{\rho}$. The proof is completed by taking the maximum over all $|\mathcal{S}_{i}|$.
@@ Appendix B Proof of Proposition 2
Our starting point is the result in Proposition 1:
@@
We bound the size of $\mathcal{S}_{i}$ using an expression that is independent of the segment index. Our bound on $|\mathcal{S}_{i}|$ is based on [9, Equation 14]:
$|\mathcal{S}_{i}|\;\;\leq$   $\;\;4\left\lfloor\frac{\rho N}{2}\right\rfloor+4\left\lfloor\frac{\rho N}{2}\right\rfloor\cdot\left\lfloor\frac{\|f(t)\|_{\infty}-\lambda^{\prime}}{2\lambda^{\prime}}\right\rfloor$
$\leq$   $\;\;2\rho N+2\rho N\cdot\left(\frac{\|f(t)\|_{\infty}-\lambda^{\prime}}{2\lambda^{\prime}}\right)$ (26)
@@
as the modulo threshold for equation (27) gives us $\mathrm{OF}\geq 3$. However, since $\lambda^{\prime}\leq\|f(t)\|_{\infty}$, then
$\mathrm{OF}\cdot\left[1-\frac{K_{\mathrm{SL}}}{N}\right]-2>1.$

### v2

@@ IV-A Analysis of the Sufficient Condition for $b$
IV-B Theoretical vs. Simulated MSE of the Sliding DFT-based Unfolding
@@
where $p_{\mathrm{rc}}(t)$ is a raised-cosine pulse with roll-off parameter $\beta=1$. The peak amplitude of the raised-cosine pulse $p_{\mathrm{rc}}(t)$ is set to unity. The pulse amplitude $A_{m}$ is drawn uniformly from the discrete set $\{-2,-1,0,+1,+2\}$, and $T$ is the time difference between two consecutive pulses. Intuitively, the signal $f(t)$ is constructed by generating 100,000 raised-cosine pulses with random amplitudes at every symbol period $T$. This signal is fed to a low-resolution 4-bit modulo ADC with sampling rate $\frac{1}{T_{\mathrm{s}}}=\frac{\mathrm{OF}\omega_{m}}{2\pi}$ to produce the quantized modulo samples $f_{\lambda^{\prime},\mathrm{q}}[n]$. The dither sequence and modulo threshold are configured according to the modulo ADC parameters. Note that it is not practical to unfold $f(t)$ using the DFT-based methods described in [10, 13, 14, 11, 36] due to its large $N_{0}$ and its lack of periodicity.
@@ V Summary and Future Directions
It is worth mentioning that while the proposed unfolding scheme relies on the availability of 1-bit folding information signal $c[n]$, the proposed recovery can be modified to work without $c[n]$. This can be done by performing robust sparse recovery of $\underline{\tilde{z}^{(i)}_{w}}[n]$ using a sliding DFT version of the LASSO-$B^{2}R^{2}$[13]. This idea, together with its performance guarantees, is considered a potential subject of future work. We also note that the sufficient conditions derived for the oversampling factor and ADC resolution are not tight. Hence, another future direction is to tighten these sufficient conditions. The modulo sampling architecture considered in this work did not take into account non-ideal modulo nonlinearities such as hysterisis and folding transients. Thus, we also intend to apply the sliding DFT recovery to generalized modulo sampling models and analyze how modulo hysterisis and folding transients affects the algorithm’s performance. Lastly, it is also of interest to apply this proposed recovery algorithm to specific low-resolution signal processing systems (e.g., communication receivers, radars, etc.).
@@ Appendix A Proof of Proposition 1
$K\geq$ $N(1-\rho)-K_{\mathrm{SL}}.$ (23)
@@
for all $i$. Here, we used $\mathrm{OF}=\frac{1}{\rho}$. The proof is completed by taking the maximum over all $|\mathcal{S}_{i}|$.
@@ Appendix B Proof of Proposition 2
We first bound the size of $\mathcal{S}_{i}$ using an expression that is independent of the segment index. To do this, we note that a real bandlimited signal $f(t)$ crosses a fixed level $\ell$ in a time interval of length $T$ at most $2f_{B}T$ times, where $f_{\mathrm{B}}$ is the bandwidth of $f(t)$. Consider the time interval that corresponds to the $i$-th segment. Since $2f_{\mathrm{B}}=\frac{\omega_{m}}{2\pi}$ and $T=NT_{\mathrm{s}}=\frac{2\pi N}{\mathrm{OF}\cdot\omega_{m}}$, the number of times that $f(t)$ crosses level $\ell$ is at most $\frac{N}{\mathrm{OF}}$. Moreover, the number of $(2\mathbb{Z}+1)\lambda^{\prime}$ folding levels in $f(t)$ is $2+2\lfloor\frac{\|f(t)\|_{\infty}-\lambda^{\prime}}{2\lambda^{\prime}}\rfloor$ as shown in [13]. Consequently, $|\mathcal{S}_{i}|$ is bounded as follows:
$|\mathcal{S}_{i}|\;\leq$ $\;\frac{N}{\mathrm{OF}}\left(2+2\bigg{\lfloor}\frac{\|f(t)\|_{\infty}-\lambda^{\prime}}{2\lambda^{\prime}}\bigg{\rfloor}\right)$
@@
[…]
Consequently, setting
$\mathrm{OF}\geq\frac{N}{N-\frac{N}{\mathrm{OF}}\left(1+\frac{\|f(t)\|_{\infty}}{\lambda^{\prime}}\right)-K_{\mathrm{SL}}}$
ensures that Proposition 1 is satisfied by transitivity. We can isolate $\mathrm{OF}$ on one side to get
@@
as the modulo threshold satisfies the inequality in (27). However, since $\lambda^{\prime}\leq\|f(t)\|_{\infty}$, then
$\mathrm{OF}\cdot\left[1-\frac{K_{\mathrm{SL}}}{N}\right]-2\geq 1.$